Hello,
Can someone please help me with this question?
What is the mathematical difference between a pH of 5.0 and a pH of roughly 4.0 - is it really that big of a difference?
Thank you for your help

Answer by drk(1908)

The pH scale is logarithmic and as a result, each whole pH value below 7 is ten times more acidic than the next higher value. For example, pH 4 is ten times more acidic than pH 5 and 100 times (10 times 10) more acidic than pH 6.
--
The concentrations of hydrogen ions and indirectly hydroxide ions are given by a pH number. pH is defined as the negative logarithm of the hydrogen ion concentration. The equation is:
pH = - log [H+]
--
Example: If an acid has an H+ concentration of 0.0001 M, find the pH.
Solution:
First convert the number to exponential notation, find the log, then solve the pH equation.
H+ = 0.0001M = 10-4; log of 10-4 = -4;
pH = - log [ H+] = - log (10-4) = - (-4) = +4 = pH
The purpose of the negative sign in the log definition is to give a
positive pH value.
--
The PH of coffee = 5
The PH of orange juice = 4